# VisualLens: Personalization through Visual History

### v1

@@ Abstract
We hypothesize that a user’s visual history with images reflecting their daily life, offers valuable insights into their interests and preferences, and can be leveraged for personalization. Among the many challenges to achieve this goal, the foremost is the diversity and noises in the visual history, containing images not necessarily related to a recommendation task, not necessarily reflecting the user’s interest, or even not necessarily preference-relevant. Existing recommendation systems either rely on task-specific user interaction logs, such as online shopping history for shopping recommendations, or focus on text signals. We propose a novel approach, VisualLens, that extracts, filters, and refines image representations, and leverages these signals for personalization. We created two new benchmarks with task-agnostic visual histories, and show that our method improves over state-of-the-art recommendations by 5-10% on Hit@3, and improves over GPT-4o by 2-5%. Our approach paves the way for personalized recommendations in scenarios where traditional methods fail.
@@ 1 Introduction
Figure 1: VisualLens leverages a user’s task-agnostic visual history to provide personalized recommendations. Our method outperforms GPT-4o by 1.6$\sim$4.6on Hit@3.
@@
Imagine a personal assistant observing what you do in your daily life. With her keen insight, she can make informed guesses about what you may enjoy or find intriguing. When you ask for recommendations on anything from restaurants and activities to movies, books, and products, based on her in-depth understanding of you she will provide suggestions tailored specifically to your tastes.
Although the idea seems straightforward, a comprehensive personal assistant that can make recommendations across various aspects of life has yet to become a reality. Most existing recommendation systems are limited to specific domains and rely on interaction history within the particular verticals (Tan and Jiang, 2023; Papadakis et al., 2023). For instance, an e-commerce platform may suggest products based on your past purchases, but it will not take into account your dining habits or interests outside of shopping. Similarly, a restaurant booking system may recommend eateries based on our previous reservations, but it will not know your new interest in cooking or your recent travels. A video streaming service may propose movies based on your viewing history, but it will not be aware of your cultural curiosities.
Recent advances in personalized conversation systems have utilized task-agnostic conversation logs to provide personalized answers (Harper and Konstan, 19; Ni et al., 2019; Liu et al., 2023). However, these approaches predominantly rely on textual data.
We hypothesize that a user’s visual history, comprising images reflecting their daily life, offers valuable insights into their interests and preferences. This work aims to investigate whether such a visual history can enable a recommendation system to better understand a user’s individual preferences, and deliver more personalized recommendations.
There are many challenges in achieving this goal. First, although we envision a future system, akin to Vannevar Bush’s Memex (Bush, 1945), that can record a user’s visual history with her consent, such a system remains elusive because of hardware limitations such as battery life, thermal constraints, and storage capacity. Second, a user’s visual history is inherently diverse and noisy, containing images not directly related to a specific recommendation ask, entities not accurately reflecting the user’s preferences (e.g., visiting a museum without enjoying it), and elements not preference-indicative (e.g., background objects like trash cans). Lastly, since many recommendation tasks require instant responses, we need to develop a low-latency solution.
In this paper, we propose a novel approach VisualLens, as a first step towards harnessing a user’s visual history for personalized recommendation. Our first contribution tackles the challenge of history availability by leveraging photos taken by a user. Unlike extensive visual logs, the photos a user takes require significantly less storage, oftentimes are readily available in reviews and social media posts, and convey more meaningful signals about a user’s interests and preferences. To facilitate evaluation, we created two benchmarks—Google Review-V and Yelp-V—using publicly available photos, providing a foundation for personalization assessments.
Our second contribution is a recommendation solution that leverages a user’s visual history to improve personalization. At the core of our solution is a set of models that effectively extract the essential signals from the visual history to infer the user’s preferences. First, given a recommendation request, we selectively retrieve only the most relevant images, thereby reducing unnecessary noises and distractions. Second, to effectively capture the signals a photo may convey, we use not only visual embeddings but also text captions and aspect words extracted from the image. Third, we adopt an iterative refinement process that continually improves aspect extraction to better reflect user interests and inform recommendations. By jointly training aspect word extraction and prediction tasks within a unified model, VisualLens not only reduces the overall parameter size, but also enhances the model’s capability to understand and utilize the visual history for accurate recommendation.
Our third contribution is a pipeline that applies multiple optimizations to accelerate recommendation generation. To minimize online processing time, we pre-generate image captions and extract aspects offline. Furthermore, we introduce a grid-based approach that aggregates multiple images into a $d\times d$ grid, enabling simultaneous processing of multiple images and significantly reducing computation overhead.
@@ 2 Related Works
Recommendation system with large language models.
@@
Large Language Models (LLMs) have demonstrated strong potential in recommendation systems with their advanced language processing capabilities (Tan and Jiang, 2023). On item-based recommendation, studies such as LLM4RS (Dai et al., 2023), LLMRank (Hou et al., 2024), CLLM4Rec (Zhu et al., 2024b), P5 (Geng et al., 2022), and Sanner et al. (2023) explored various LLM prompt and bootstrapping strategies, showing competitive performance, especially in cold-start scenarios. Generative recommendation with open-domain prompts is explored by GenRec (Ji et al., 2023), though fine-tuning remains crucial. Fine-tuning approaches include personalized aspect extraction (Li et al., 2023c), and multitasking on LLaMa models (Yang et al., 2023). Retrieval-enhanced models, such as ReLLa, improve recommendation by retrieving relevant behavior sequences (Lin et al., 2024). Instruction-tuning and graph augmentation approaches are explored in InstructRec (Zhang et al., 2023), LLMRec (Wei et al., 2024b), and LKPNR (Hao et al., 2023). LLMs also work well with content-based recommendation, such as Li et al. (2023a) on prompting, and ONCE (Liu et al., 2023), LoID (Xu et al., 2024) on fine-tuning.
Multimodal large language models.
Multimodal LLMs are becoming increasingly powerful in processing images and videos and in generating human-like text, exemplified by models like GPT-4 family (OpenAI, 2024), Claude 3.5 family (Anthropic, 2024), Gemini and PaliGemma (Google, 2023; Beyer et al., 2024), LLaVA model family (Liu et al., 2024), and Llama 3 Vision models (Meta, 2024). However, they still suffer from a strong language prior (Fu et al., 2024a; Parcalabescu and Frank, 2024), generalization (Zhu et al., 2022) and hallucinations (Fu et al., 2024b), or require heavy text transcription (Zhu et al., 2024a). The extent to which they can understand user history, in analogy to LLM recommendation systems, is still unclear.
Multimodal recommendation systems.
@@
Multimodal recommendation systems leverage multiple data types, such as text and images, to improve recommendation relevance and personalization. Before the LLM era, Lee and Abu-El-Haija (2017) proposed systems for content-only video recommendation using similarity learning. PC2(Yan et al., 2023) develop an LLM model that provides multimodal explanations for recommendations. On modalities beyond image and text, LSVCR (Zheng et al., 2024) built a joint recommendation system integrating comments with video items (Davidson et al., 2010; Du et al., 2020; Gao et al., 2017; Zhou et al., 2024). The current state-of-the-art image-text recommendation, UniMP (Wei et al., 2024a), extended single-task multimodal personalization (He and McAuley, 2016; Tang et al., 2020; Wei et al., 2019, 2023) on multitask website-based shopping. Methods such as RecFormer (Li et al., 2023b) and UniSRec (Hou et al., 2022) convert images to short captions to utilize text-only models. However, none of these methods consider the scenario of personalized recommendation from the user’s visual history. In this work, we present a novel method VisualLens leveraging multimodal LLM to achieve personalization with visual history.
Figure 2: VisualLens inference pipeline: the offline process augments images in the visual history with captions and aspect words; the runtime recommendation process retrieves relevant images, generate user profile accordingly, and then predict candidate preferences.
3 Problem Definition
@@
Consider a recommendation QA (question answering) task, where the user asks a recommendation question $q$, and the recommender answers $q$ with a ranked list of items. Good recommenders shall rank the items that the user is more likely to be interested in or willing to try early in the list.
@@
The recommender is facilitated with a task-agnostic visual history ${\cal H}_{u}$ for each user $u$, which contains a series of photos, taken or posted by the user, not necessarily related to $q$. We state three assumptions to allow generalization. First, the photos may not be directly relevant to the question. Second, an image may not be associated with any candidate item, and even so, the candidate ID is not given. Third, a photo does not necessarily imply strong preferences. Figure 1 shows an example question, visual history, and candidates.
@@
To simplify the problem, we assume a candidate retriever exists to retrieve all candidates that satisfy the user’s question. Each candidate $s$ is represented with a $(x_{s},{\cal I}_{s})$ pair, where $x_{s}$ is the name and text descriptions, and ${\mathcal{I}}_{s}$ is the image set for the item.
@@
Traditional recommendation setting considers two types of signals. The first is a task-specific set of items in the candidate set, which captures user interest or at least user history. The second is a set of user-specific attributes such as the user’s age, gender, and interests. This paper focuses on task-agnostic visual history and leaves integration of these traditional signals for future extensions.
4 Recommendation Method
We start with describing the VisualLens solutions. Our solution contains two parts: offline history augmentation, and runtime recommendation generation.
  • Offline history augmentation4.1) augments each image in the visual history with captions and aspect words extracted from the image. Figure 2 shows example captions and aspect words of the images in the visual history.
@@ 4 Recommendation Method
  • Runtime recommendation generation4.2) answers a recommendation question $q$ in three steps. First, the history retrieval step retrieves only images relevant to $q$, since the visual history can be diverse and not all photos are relevant to every query. Second, the preference profiling step uses the retrieved images and their augmented captions and aspects to generate the user’s preference profile. Third, the candidate matching step matches the user’s preference profile with each candidate, to generate the confidence score for each candidate for ranking. VisualLens utilizes the same backbone model with parameter $\theta$ for aspect generation and candidate matching.
4.1 Offline history augmentation
To best capture various signals in an image, we encode it to visual embeddings and augment it with a caption and a set of aspect words. As we show in the experimental ablation study (Table 6.3), these augmentations allow us to leverage the strengths of various encoding techniques and considerably improve the recommendation quality. These augmentations can be used directly at recommendation time to reduce runtime latency.
Image encoding.
We encode each image in the visual history using the CLIP ViT-L/14@336px model (Radford et al., 2021). This embedding will be directly leveraged in image retrieval in the history retrieval step at runtime.
Caption generation.
We generate image captions using a frozen LLaVA-v1.6 8B model (Liu et al., 2024). We prompted the model to generate captions within 30 words, such that the captions are succinct and short of hallucinations.
Aspect word generation.
Aspect words are descriptive terms that capture essential features or attributes of an image (e.g., dome, balcony, plants). Aspect words provide crucial details about the image. We prompted the backbone model to list aspect words for an image. We do not constrain the number of extracted aspect words since each image may have varying numbers of aspect words essential for final preference prediction. We will describe in Section 5.3 how we finetune the model to generate aspect words that are most useful for recommendation tasks.
4.2 Runtime recommendation generation
History retrieval.
@@ 4.2 Runtime recommendation generation
Given a query $q$ and a user’s visual history ${\cal H}_{u}$, VisualLens first retrieves images that are related to $q$, denoted by $\mathcal{I}_{u,q}$. We choose up to $w$ images to cap the number of images we process at runtime, and ensure that only the most contextually relevant images are retained for further processing, thereby reducing noise.
@@
In general, we can use any image retrieval method such as DELG (Cao et al., 2020). Here, we present a method for categorical recommendations such as restaurants and museums, popular in recommendation tasks. For each category $c$, we randomly select a set of candidate items in the category and average the visual embeddings of their images as the image embedding of the category, denoted by $\mathbf{v}_{c}$. Specifically, the category embedding is calculated as:
$\mathbf{v}_{c}=\frac{1}{n}\sum_{j=1}^{n}\mathbf{v}_{c}^{(j)}$
where $n$ is the number of candidates, and $\mathbf{v}_{c}^{(j)}$ indicates the visual embedding of the $j$-th item image in category $c$. The retrieval step measures the cosine similarity between the visual embedding $\mathbf{v}_{i}$ of each image $i\in{\cal H}$ in the user’s history and the image embedding $\mathbf{v}_{c}$ of the relevant category $c$. We then select the top-$w$ images based on the cosine similarity scores.
Preference profiling
. Given a set of retrieved images $\mathcal{I}_{u,q}$, VisualLens then generates the user’s preference profile relevant to the images, their captions, and aspect words.
@@
A critical part of this step is image encoding. Even after retrieval, the number of images $w$ is still large. Most multimodal LLMs allow context windows of limited sizes, constraining the number of images we can process. For example, for an input image of resolution $896\times 896$, the PaliGemma model would generate an embedding of up to 4,096 tokens. A typical LLM with a context window of 8,192 tokens can take at most 2 images.
@@
We propose to group relevant images $\mathcal{I}_{u,q}$ into a $d\times d$ grid, where $d^{2}=w$, and treat all images in the grid as a single image. If we have retrieved fewer than $w$ images, we pad with a black background. Let $h$ be the maximum available resolution in a multimodal LLM. The gridify process $G$ takes the $d\times d$ grid and generates an image of fixed size $\mathbb{R}^{h\times h\times 3}$. Additionally, we number each image from 1 to $d^{2}$ to ensure the images are grounded to the corresponding caption and aspect words in the input to the candidate matching.
We denote a user $u$’s preference profile on question $q$ by $(i_{u,q},x_{u,q})$, where $i_{u,q}$ denotes the gridified image, and $x_{u,q}$ denotes the concatenated captions and aspect words of relevant images.
Candidate matching.
@@
Finally, VisualLens takes the user’s preference profile $(i_{u,q},x_{u,q})$ and a set of candidates, each represented by $(x_{s},{\mathcal{I}}_{s})$, and generates the matching score for each candidate, which will then be used for ranking. This is achieved by prompting the multimodal candidate predictor, where we packed the user profile and candidates to the prompt through the image channel and the text channel separately (see the template in Appendix C).
5 Iterative Refinement and Joint Training
VisualLens requires LLMs for image encoding, image captioning, aspect word generation, and final preference-candidate matching. Out-of-box LLMs do not perform well, and we conduct fine-tuning in three steps.
  • First, to better encode the $d\times d$ grid of images, we fine-tune image encoding with multi-image caption tasks.
  • Second, since aspect generation is not a standard multimodal task, we iteratively refine the aspect word extractor, ensuring that the keywords precisely capture the aspects of the image in a suitable granularity.
  • Third, to best leverage personal profile signals including image embeddings, captions, and aspect words, we finetune the candidate matching model. This is done through joint multitask training with the aspect words generation.
5.1 Multi-image caption pretraining

### v2

# VisualLens: Personalization through Task-Agnostic Visual History

@@ Abstract
Existing recommendation systems either rely on user interaction logs, such as online shopping history for shopping recommendations, or focus on text signals. However, item-based histories are not always accessible, and are not generalizable for multimodal recommendation. We hypothesize that a user’s visual history — comprising images from daily life — can offer rich, task-agnostic insights into their interests and preferences, and thus be leveraged for effective personalization. To this end, we propose VisualLens, a novel framework that leverages multimodal large language models (MLLMs) to enable personalization using task-agnostic visual history. VisualLens extracts, filters, and refines a spectrum user profile from the visual history to support personalized recommendation. We created two new benchmarks, Google Review-V and Yelp-V, with task-agnostic visual histories, and show that VisualLens improves over state-of-the-art item-based multimodal recommendations by 5-10% on Hit@3, and outperforms GPT-4o by 2-5%. Further analysis shows that VisualLens is robust across varying history lengths and excels at adapting to both longer histories and unseen content categories.
@@ 1 Introduction
Imagine a personal assistant, similar to Vannevar Bush’s Memex (Bush, 1945), observing what you do in your daily life. With her keen insight, she can make informed guesses about what you may enjoy or find intriguing. When you ask for recommendations on anything from restaurants and activities to movies, books, and products, based on her in-depth understanding of you she will provide suggestions tailored specifically to your tastes.
While the concept is intuitive, a truly comprehensive personal assistant capable of making recommendations across all aspects of life has yet to be realized. Most existing multimodal recommendation systems remain domain-specific and rely heavily on item-based interaction histories (Tan and Jiang, 2023; Papadakis et al., 2023). For example, an e-commerce platform may suggest products based on past purchases but ignore dining habits or interests outside shopping.
This work explores how can we leverage such a user’s visual record to better understand individual preferences and enable more general, personalized recommendations. Achieving task-agnostic recommendations from visual history poses several challenges. First, visual histories are often diverse and noisy, containing images unrelated to any specific recommendation task, entities that fail to accurately represent user preference, or non-informative elements (e.g., background objects like trash cans). This creates a trade-off between preserving rich visual content and extracting clean, interpretable representations of user interests. Second, most current MLLMs can only process a limited number of images, requiring selective retrieval of relevant user profile information for each query. Third, existing benchmarks are inadequate for evaluating task-agnostic visual recommendation systems, highlighting the need for new, purpose-built evaluation datasets.
We propose a novel framework VisualLens, as a first step towards harnessing a user’s visual history for MLLM recommendation. VisualLens begins by extracting an offline spectrum user profile, compressing each image in the visual history into a triplet: (raw image, caption, aspect words). This representation spans a spectrum from rich but noisy content (raw image) to concise but clean semantic cues (aspect words). To improve the quality of aspect words, we employ an iterative refinement process that progressively enhances their alignment with user interests. Next, to efficiently incorporate multiple images during runtime recommendation, VisualLens retrieves the most relevant segments of the user profile based on the query context. These selected images are organized into a $d\times d$ visual grid, accompanied by their corresponding captions and aspect words, enabling the model to jointly process and reason over a compact yet informative representation of the user’s visual history. Finally, we train the system to perform both aspect refinement and recommendation question-answering (QA) in a unified model. This design not only reduces parameter overhead but also strengthens the model’s ability to interpret and utilize visual history for accurate, personalized recommendations.
To facilitate the evaluation, we created two new benchmarks, Google Review-V and Yelp-V, providing a foundation for personalization assessments. We leverage user-taken photos to address the challenge of history availability. Unlike extensive Memex video logs, these photos require less storage, often available in reviews and social media posts, and provide more insights into a user’s interests and preferences. Each benchmark includes a standard test set targeting generalization to new users, along with two additional test sets for transferring to longer histories and unseen categories.
Our experimental study shows promising recommendation quality of VisualLens. It achieved 82-91% Hit@10 on the Google Review-V and Yelp-V benchmarks, outperforming state-of-the-art (UniMP Wei et al. (2024a)) by $\sim$$10\%$. Even comparing with GPT-4o, our 8B model improves Hit@3 by 1.6% and 4.6% respectively on the two benchmarks. Further analysis reveals that VisualLens excels in adapting to longer histories and unseen categories, while maintaining robustness with shorter histories.
@@ 2 Related Works
Recommendation system with large language models. Large Language Models (LLMs) have demonstrated strong potential in recommendation systems with their advanced language processing capabilities (Tan and Jiang, 2023). On item-based recommendation, studies such as LLM4RS (Dai et al., 2023), LLMRank (Hou et al., 2024), CLLM4Rec (Zhu et al., 2024b), P5 (Geng et al., 2022), and Sanner et al. (2023) explored various LLM prompt and bootstrapping strategies, showing competitive performance, especially in cold-start scenarios. Generative recommendation with open-domain prompts is explored by GenRec (Ji et al., 2023), though fine-tuning remains crucial. Fine-tuning approaches include personalized aspect extraction (Li et al., 2023c), and multitasking on LLaMa models (Yang et al., 2023). Retrieval-enhanced models, such as ReLLa, improve recommendation by retrieving relevant behavior sequences (Lin et al., 2024). Instruction-tuning and graph augmentation approaches are explored in InstructRec (Zhang et al., 2023), LLMRec (Wei et al., 2024b), and LKPNR (Hao et al., 2023). Recent advances in personalized conversation systems have utilized task-agnostic conversation logs to provide personalized answers (Li et al., 2023a; Harper and Konstan, 19; Ni et al., 2019; Liu et al., 2023; Xu et al., 2024). However, these content-based recommendation approaches predominantly rely on textual data.
@@
Multimodal recommendation systems. Multimodal recommendation systems (Wang et al., 2022; Yang et al., 2024; Sheng et al., 2024) leverage multiple data types, such as text and images, to improve recommendation relevance and personalization. Before the LLM era, Lee and Abu-El-Haija (2017) proposed systems for content-only video recommendation using similarity learning. PC2L (Yan et al., 2023) develop an LLM model that provides multimodal explanations for recommendations. On modalities beyond image and text, MMRF (Zhou et al., 2024) built a joint recommendation system integrating comments with video items (Davidson et al., 2010; Du et al., 2020; Gao et al., 2017; Zheng et al., 2024). RecFormer (Li et al., 2023b) and UniSRec (Hou et al., 2022) convert images to short captions to utilize text-only models. The current state-of-the-art image-text recommendation, UniMP (Wei et al., 2024a), extended single-task multimodal personalization (He and McAuley, 2016; Tang et al., 2020; Wei et al., 2019, 2023) on multitask website-based shopping. However, most existing multimodal recommendation approaches rely on item-based user history, which is not always available. To address this limitation, we propose VisualLens, a novel framework that leverages MLLMs to enable personalization using task-agnostic visual history.
Table 1: Unlike representative related works, VisualLens is a novel framework leveraging multimodal LLM for recommendation with task-agnostic visual content. CTR: click-through rate.
Multimodal New Eval User History User Profile Objective
LLM4RS (Dai et al., 2023) Textual item features Text sequence Items
ReLLa (Lin et al., 2024) Textual item features Top-k behaviors CTR
ONCE (Liu et al., 2023) Task-agnostic text Content embedding Items
PC2L (Yan et al., 2023) Multimodal item features Selected images Explanation
COURIER (Yang et al., 2024) Multimodal item features Joint embedding CTR
UniMP (Wei et al., 2024a) Multimodal item features Joint embedding Multi-task
VisualLens Task-agnostic images Spectrum QA
3 Multimodal Task-Agnostic Recommendation
@@
Consider a recommendation QA task, where the user asks a recommendation question $q$, and the recommender answers $q$ with a ranked list of candidate items. Good recommenders shall rank the items that the user is more likely to be interested in or willing to try early in the list.
@@
In multimodal task-agnostic recommendation, the recommender is facilitated with a task-agnostic visual history ${\cal H}_{u}$ for each user $u$, which contains a series of photos, taken or posted by the user, not necessarily related to $q$. We state three assumptions to allow generalization. First, the photos may not be directly relevant to the question. Second, an image may not be associated with any candidate item, and even so, the candidate ID is not given. Third, a photo does not necessarily imply strong preferences. Figure 1 shows an example question, visual history, and candidates.
@@
To simplify the problem, we assume a candidate retriever exists to retrieve all candidates that satisfy the user’s question. Each candidate $s$ is represented with a $(x_{s},{\mathcal{I}}_{s})$ pair, where $x_{s}$ is the name and text descriptions, and ${\mathcal{I}}_{s}$ is an optional image set for the item.
@@
Traditional recommendation setting considers two more types of signals. The first is a task-specific set of items in the candidate set, which captures user interest or at least user history. The second is a set of user-specific attributes such as the user’s age, gender, and interests. This paper focuses on task-agnostic visual history and leaves integration of these traditional signals for future extensions.
Figure 2: VisualLens inference pipeline: the offline process augments images with captions and aspect words to generate a spectrum user profile; the runtime recommendation process retrieves relevant images, generate query-specific user profile accordingly, and then predict candidate preferences.
4 VisualLens Framework
@@
VisualLens framework contains two parts in inference (Figure 2): offline user profile generation, and runtime recommendation. Offline user profile generation (§4.1) augments each image in the visual history with captions and aspect words, and builds a spectrum user profile. Runtime recommendation (§4.2) answers a recommendation query $q$ in three steps. First, the history retrieval step retrieves only images relevant to $q$, since the visual history can be diverse and not all photos are relevant to every query. Second, the preference profiling step uses the retrieved images and their augmented captions and aspects to generate a query-specific profile of the user. Third, the candidate matching step matches the query-specific profile with each candidate, to generate the confidence score for each candidate for ranking. We discuss the joint training algorithm of VisualLens in §4.3.
4.1 Offline user profile generation
To build a user profile that retains rich visual content while offering clean, interpretable cues about user interests, we augment each image with a caption and a set of aspect words. Each image is thus represented as a spectrum triplet — (raw image, caption, aspect words) — ordered by decreasing information richness and increasing semantic clarity. Ablation results (Table 5.3) confirm that these augmentations improve recommendation performance.
Image encoding. Each image is encoded using the CLIP ViT-L/14@336px model (Radford et al., 2021), producing embeddings used for history retrieval at query time.
Caption generation. Captions are generated using a frozen LLaVA-v1.6-8B model (Liu et al., 2024), prompted to produce concise ($\leq$30 words) and grounded descriptions to minimize hallucinations.
Aspect word generation. Aspect words are concise descriptors of key image attributes (e.g., dome, balcony). We prompt the model to list relevant terms without constraining the count, allowing flexibility based on image complexity.
All modules are plug-and-play and can be replaced with stronger alternatives. While LLaVA-v1.6 occasionally produces irrelevant or generic aspect words (e.g., blue, sky), we describe in Section 4.3.3 how joint finetuning improves their utility for recommendation tasks.
4.2 Runtime recommendation
@@
History retrieval. Given a query $q$ and a user’s visual history ${\cal H}_{u}$, VisualLens first retrieves images that are related to $q$, denoted by $\mathcal{I}_{u,q}$. We choose up to $w$ images to cap the number of images we process at runtime, and ensure that only the most contextually relevant images are retained for further processing, thereby reducing noise.
@@
In general, we can use any image retrieval method such as DELG (Cao et al., 2020). Here, we present a method for categorical recommendations such as restaurants and museums, popular in recommendation tasks. For each category $c$, we randomly select a set of candidate items in the category and average the visual embeddings of their images as the image embedding of the category, denoted by $\mathbf{v}_{c}$. Specifically, the category embedding is calculated as $\mathbf{v}_{c}=\frac{1}{n}\sum_{j=1}^{n}\mathbf{v}_{c}^{(j)}$, where $n$ is the number of candidates (see Appendix C for sensitivity), and $\mathbf{v}_{c}^{(j)}$ indicates the visual embedding of the $j$-th item image in category $c$. The retrieval step measures the cosine similarity between the visual embedding $\mathbf{v}_{i}$ of each image $i\in{\cal H}$ in the user’s history and the image embedding $\mathbf{v}_{c}$ of the relevant category $c$. We then select the top-$w$ images based on the cosine similarity scores.
@@
Preference profiling. Given a set of retrieved images $\mathcal{I}_{u,q}$, VisualLens then generates user’s query-specific profile. A critical part of this step is image encoding. Even after retrieval, the number of images $w$ is still large. Most MLLMs allow context windows of limited sizes, constraining the number of images we can process. For example, for an input image of resolution $896\times 896$, the PaliGemma model would generate an embedding of up to 4,096 tokens. A typical LLM with a context window of 8,192 tokens can take at most 2 images.
@@
We propose to group relevant images $\mathcal{I}_{u,q}$ into a $d\times d$ grid, where $d^{2}=w$, and treat all images in the grid as a single image. If we have retrieved fewer than $w$ images, we pad with a black background. Let $h$ be the maximum available resolution in a multimodal LLM. The gridify process $G$ takes the $d\times d$ grid and generates an image of fixed size $\mathbb{R}^{h\times h\times 3}$. Additionally, we number each image from 1 to $d^{2}$ to ensure the images are grounded to the corresponding caption and aspect words in the input to the candidate matching. We denote a user $u$’s profile on question $q$ by $(i_{u,q},x_{u,q})$, where $i_{u,q}$ denotes the gridified image, and $x_{u,q}$ denotes the concatenated captions and aspect words of relevant images.
@@
Candidate matching. Finally, VisualLens takes the query-specific user profile $(i_{u,q},x_{u,q})$ and a set of candidates, each represented by $(x_{s},{\mathcal{I}}_{s})$, and generates the matching score for each candidate, which will then be used for ranking. This is achieved by prompting the multimodal candidate predictor, where we packed the user profile and candidates to the prompt through the image channel and the text channel separately (see the prompt template in Appendix D).
Table 2: Dataset statistics of Google Review-V and Yelp-V.
Dataset Train Dev Test Categories Avg. # of images Avg. # of GT Avg. # of candidates
GR-V 15.69M 2K 200K 66 157.0 2.7 43.1
Yelp-V 4.12M 2K 100K 35 263.6 8.2 66.7
4.3 Iterative Refinement and Joint Training
VisualLens relies on LLMs for image encoding, caption generation, aspect word extraction, and profile–candidate matching. Current off-the-shelf models perform suboptimally for these tasks, so we apply continued pretraining and task-specific fine-tuning to enhance performance.
4.3.1 Multi-image caption pretraining